# Implementing Evidence Acquisition:

### v2

@@ 6.1 Randomized Contracts
Our goal has been to determine the maximum effort implementable within the class of contracts defined in Section 2.2. As histories only include messages sent by the agent, we implicitly rule out the use of randomization (as in, for instance, Deb et al., (2018)). Deterministic mechanisms have significant practical appeal, as it is not always obvious what a randomization might correspond to and how a mechanism designer might implement this. These issues have been discussed extensively in the contracting literature; we refer the reader to discussions in Laffont and Martimort, (2002) as well as Bester and Strausz, (2001) on this point to avoid detours. Now, randomization provides no additional benefit in implementing maximum effort with probability 1. But it is natural to ask whether some designer objectives may yield benefits to randomization. Our analysis speaks to this question as well.
@@
Crucially, in randomized contracts, at any time $t$, the history of the randomization device $\left\{\varsigma_{t^{\prime}}\right\}_{t^{\prime}\leq t}$ is publicly revealed to the agent before determining his choice of effort or the message sent to the designer. The randomization revealed before $t$ affects the agent’s incentives after time $t$, and without it, such contracts reduce back to deterministic contracts.
@@
Nevertheless, our analysis suggests randomization can expand the set of implementable strategies. We previously observed that it may be possible to get the agent to work from $\mu_{1}$ to $\mu_{2}$ and to get the agent to work from $\mu_{2}$ to $\mu_{3}$, but not from $\mu_{1}$ to $\mu_{3}$. This would occur if the reward necessary to get the agent to work to $\mu_{3}$ were so high that the agent would “shirk-and-lie” at $\mu_{1}$. However, the agent may be willing to start working at $\mu_{1}$, not knowing whether the reward will be “high” or “low”—but once the agent starts working, the designer can randomly inflate or decrease the rewards of the agent. Once time has passed, the outcome of the randomization can be revealed, and if the rewards inflate, the agent has incentives to exert effort for longer absent a Poisson signal arrival—so that the realized stopping time increases for some realization of the randomization.
@@
We illustrate this intuition more formally and rigorously for the environment from Section 3.1, where learning is perfect and only good news signal $G$ arrives with positive probability. As discussed, under deterministic contracts, a V-shaped scoring rule with parameters $r_{0}=1$ and $r_{1}\in[0,1]$ implements the effort-maximizing contract $R$; in particular, Section 3.1 characterized when in fact effort-maximizing contracts require $r_{1}<1$. Recall that we denote $\tau_{R}$ as the stopping time in the effort-maximizing deterministic contract and let $\mu^{N}_{\tau_{R}}$ be the stopping belief when no Poisson signal is observed. Let $\delta\in(0,1-r_{1}]$ be the maximum number such that (1) the agent has strict incentives to exert effort until time $\frac{2\tau_{R}}{3}$ absent signal arrival given menu options $(1,0)$ and $(0,r_{1}-\delta)$; and (2) the agent can be incentivized to exert effort given menu options $(1,0)$ and $(0,r_{1}+\delta)$ given belief $\mu^{N}_{\frac{\tau_{R}}{2}}$. Now consider the randomized contract $\hat{R}$ that provides menu options $(1,0)$ and $(0,r_{1}-\delta)$ from time 00 to $\frac{\tau_{R}}{2}$, and after time $\frac{\tau_{R}}{2}$, offers menu options $(1,0)$ and $(0,r_{1}+\delta)$ with probability $\epsilon^{2}$, offers menu options $(1,0)$ and $(0,0)$ with probability $\epsilon$, and offers the same menu options $(1,0)$ and $(0,r_{1}-\delta)$ otherwise. With sufficiently small $\epsilon>0$, the agent still has incentives to exert effort at any time $t\leq\frac{\tau_{R}}{2}$. Moreover, after time $\frac{\tau_{R}}{2}$, with probability $\epsilon^{2}$, the realized menu options are $(1,0)$ and $(0,r_{1}+\delta)$, and the agent can be incentivized to exert effort to a time strictly larger than $\tau_{R}$ in the absence of a Poisson signal. [^12] This kind of modification may be of interest to a designer who only values extreme posterior beliefs. For instance, suppose the designer faced a decision problem where the possible decisions belonged to a set $A=\{0,1\}$. Consider a designer payoff function of $v(0,\theta)=0$ for all $\theta\in\Theta$, and $v(1,0)=1,v(1,1)=-\frac{1-\mu^{N}_{\tau_{R}}}{\mu^{N}_{\tau_{R}}}$; plainly, the designer only seeks to change her action from 1 to 0 if the posterior belief is below $\mu^{N}_{\tau_{R}}$. In this case, the payoff under any deterministic contract is 0. However, under the randomized contract outlined, the designer would obtain a positive payoff when implementing the identified randomized contract.
@@ 6.2 Non-invariant Environments
6.3 Valuing Transfers and Ex-Ante Reward Constraints
As our analysis focused entirely on the problem of implementing maximum effort, we have (a) avoided degeneracy by imposing bounds in the rewards to the agent and (b) not considered the ex-ante expected value of transfers (as stated in the introduction, either viewing them as non-monetary or zero-sum). An alternative to part (a) would be to impose an ex-ante reward constraint on the designer rather than ex-post as in our main model; solving this problem for any bound allows the designer to identify tradeoffs between expected payments and information acquisition, thus addressing part (b).
Even if there is an ex-ante upper bound on the payments the agent can be provided, limited liability is still necessary to avoid degeneracy—i.e., the reward should satisfy $r\geq 0$ ex-post. To implement an arbitrary stopping strategy $\tau$, simply find a decision problem where exerting effort until time $\tau$ is optimal. If the only constraint is that payments should have expectation zero, then $\tau$ could be implemented using a “selling-the-firm-to-the-agent” mechanism—that is, charging the agent the ex-ante expected utility from that decision problem up front and then providing a reward (or punishment) equal to the ex-post utility under the decision problem. Thus, we discuss issues under this modification by focusing on the model with ex-ante reward constraints and ex-post limited liability constraints.
We show that in the special case where the conditions in both perfect-learning environments and single-signal environments are satisfied, i.e., when $\lambda^{B}_{0}=\lambda^{B}_{1}=\lambda^{G}_{0}=0$ and only $\lambda^{G}_{1}>0$, the effort-maximizing contract can also be implemented as a scoring rule. Appendix OA 5 contains details for this claim. [^14] This result follows immediately from Proposition 1 of Gerardi and Maestri, (2012), but we present an independent proof for completeness. This observation illustrates that the optimality of static contracts extends under ex-ante reward constraints under natural assumptions. Continuing with the above, these results also imply that when only $\lambda^{G}_{1}>0$, the value of ex-ante expected payments implementing some effort level $\tau$ cannot be minimized by resorting to dynamic contracts. We conjecture that our result also extends when only conditions in one of our canonical environments are satisfied (rather than both), although the formal arguments for this conjecture may require additional novel ideas.
7 Conclusion
@@
We have articulated how dynamic rewards can expand the set of implementable strategies in a simple yet fundamentally dynamic information acquisition problem. The economic importance of contracting for information acquisition is self-evident, and most natural stories for why information acquisition is costly involve some dynamic element. Our goal has been to take such dynamics seriously, for learning technologies with a natural interpretation in terms of investigators seeking evidence, and under a class of contracts reflecting the provision of bonuses for correct advice.
@@
As our focus is on contracting under a general class of mechanisms, a fundamental difficulty underlying our exercise is the lack of any natural structure (e.g., stationarity) under an arbitrary dynamic contract. Such assumptions are often critical in similar settings. Despite this fundamental challenge, we provided simple, economically meaningful conditions such that maximum effort is implementable by a scoring rule and explained the extent to which these conditions are necessary for this conclusion to hold.
@@
There are many natural avenues for future work. Empirically, our results show how learning technologies influence the benefits of time variation to bonuses. A natural question is whether considerations our model has not captured influence the extent to which contracts do or do not involve such features. Theoretically, accommodating both dynamic moral hazard and adverse selection required us to focus on a class of technologies with recognized significance but less generality than, for instance, Chambers and Lambert, (2021). Still, we do not doubt similar conclusions could emerge under different information arrival processes. More broadly, we view questions regarding whether or not simple contracts are limited in power relative to dynamic mechanisms as a worthwhile agenda overall. Any further insights on these questions would prove valuable toward understanding how dynamics influence mechanism design for information acquisition, for both theory and practice.
@@ A.1 Menu Representation
A.2 Details Behind Section 3.1
Here we provide some additional details behind the calculations in Section 3.1. We consider any scoring rule which involves the choice between $(r_{0},0)$ and $(0,r_{1})$, with the former being chosen when stopping in the absence of any signal arrival and the latter being chosen if one does occur. We note that stopping and accepting contract $(r_{0},0)$ delivers payoff $r_{0}(1-\mu_{t}^{N})$; if, at time $\tau$, the agent continues for a length $\Delta$ and then stops, the payoff is:
Imposing indifference between stopping and continuing yields:
From this, we obtain the following differential equation:
@@ 2. Case 2:
Online Appendix for “Implementing Evidence Acquisition: Time Dependence in Contracts for Advice”
@@
Appendix OA 2 Efficient Algorithms for Computing Effort-Maximizing Contracts
Let $\tau\leq T$ be the stopping time in the effort-maximizing contract. We know that there exists a contract $R$ such that the agent has incentive to exert effort in all continuation $\mathcal{G}_{t,\tau}$ with $t\leq\tau$. Therefore, to compute the effort-maximizing contract, it is sufficient to enumerate all stopping time $\tau$ and verify if there exists such a contract for incentivizing the agent to work until $\tau$. The effort-maximizing contract is the contract for the largest $\tau$.
To show this, fixing a stopping time $\tau$, we consider the menu representation in Lemma 2, i.e, a sequence of menu options $\{r^{s}_{t}\}_{t\leq\tau,s\in S}\cup\{r^{N}_{\tau}\}$. Moreover, we construct additional variables $\{r^{N}_{t}\}_{t<\tau}$ to simplify the constraints we need in the optimization problem. For any $t\leq t^{\prime}\leq\tau$ and $s\in S$, let $f^{s}_{t}(t^{\prime})$ be the probability of receiving Poisson signal $s$ at time $t^{\prime}$ conditional on not receiving Poisson signals before time $t$. Let $F^{s}_{t}(t^{\prime})$ be the corresponding cumulative probability. The constraints for providing sufficient incentives for the agent to exert effort is
$\left(1-\sum_{s\in S}F^{s}_{t}(\tau)\right)\cdot u(\mu^{N}_{\tau},r^{N}_{\tau})+\sum_{t^{\prime}=t}^{\tau}\sum_{s\in S}u(\mu^{s}_{t^{\prime}},r^{s}_{t^{\prime}})\cdot f^{s}_{t}(t^{\prime})\geq u(\mu_{t-\Delta}^{N},r_{t-\Delta}^{N})$
for all $t\leq\tau$ and $s\in S$. Moreover, the contract $R$ is incentive compatible if
$u(\mu_{t}^{s},r_{t}^{s})\geq u(\mu_{t}^{s},r_{t^{\prime}}^{s^{\prime}})$
for all $t\leq t^{\prime}\leq\tau$ and $s,s^{\prime}\in S\cup\{N\}$. Note that since the utility function $u(\mu,\cdot)$ is a linear function, in the discrete time model, the above set of constraints is a finite set of linear constraints on the menu options. This implies that whether a solution exists and finding a solution if it exists can be computed in polynomial time, and hence the effort-maximizing contract can also be found in polynomial time.
Appendix OA 3 Additional Review of Scoring Rules
@@
Appendix OA 4 Comparative Statics of Scoring Rules
@@
Appendix OA 5 Ex Ante Reward Constraints
Proposition 1.
In environments where both perfect-learning and single-signal conditions are satisfied, and the designer’s reward constraint is imposed in ex ante, there exists an effort-maximizing contract that can be implemented as a V-shaped scoring rule.
Proof.
Note that even with ex ante reward constraints, the menu representation of the effort-maximizing contracts still applies. Since only good news signal $G$ arrives with positive probability, it is sufficient to consider effort-maximizing contract $R$ with menu representation $\{r^{G}_{t}\}_{t\leq\tau_{R}}\cup\{r^{N}_{\tau_{R}}\}$.
First note that it is without loss to assume that $r^{N}_{\tau_{R},1}=0$ and $r^{G}_{t,0}=0$ for all $t\leq\tau_{R}$. The latter holds because signals are perfectly revealing and hence lowering the reward for state 00 upon receiving good news signals does not affect the agent’s incentive or the ex ante reward constraint of the designer. The reason why $r^{N}_{\tau_{R},1}=0$ is because we can always increase the reward $r^{N}_{\tau_{R},0}$ and decrease the reward $r^{N}_{\tau_{R},1}$ such that the agent’s expected reward given belief $\mu^{N}_{\tau_{R}}$ remains unchanged, and the dynamic incentives of the agent is not affected under single-signal environments. In this case, the dynamic incentives of the contract implies that $r^{G}_{t,1}$ is weakly decreasing over time.
Now consider another contract $\hat{R}$ that only offers menu options $r^{N}_{\tau_{R}}$ and $\hat{r}=(0,\hat{r}_{1})$ at all time $t$ with parameter $\hat{r}_{1}\leq r^{G}_{\tau_{R},1}$ chosen such that the agent’s incentive for exerting effort is binding at time $\tau_{R}$. Note that contract $\hat{R}$ can be implemented as a V-shaped scoring rule. We will show that $\tau_{\hat{R}}=\tau_{R}$ and the expected reward of the agent given contract $\hat{R}$ is lower, which implies that contract $\hat{R}$ is also effort-maximizing under the ex ante reward constraint.
Let $\hat{t}$ be the maximum time such that $\mu^{N}_{\hat{t}-\Delta}$ weakly prefers menu option $\hat{r}$ over $r^{N}_{\tau_{R}}$. Note that given contract $\hat{R}$, by the Envelope Theorem, the agent’s utility for exerting effort optimally is convex in his current belief with derivative larger than $-r^{N}_{\tau_{R},0}$. Therefore, for any time $t\in(\hat{t},\tau_{R}]$, the agent always has incentive to exert effort given contract $\hat{R}$. For any time $t\leq\hat{t}$, using the identical argument in the proof of Theorem 2 for the perfect-learning environments, by decreasing the rewards in menu options $r^{G}_{t}$, the agent’s decrease in utility for exerting effort in continuation game $\mathcal{G}_{t}$ is less than his decrease in utility for not exerting effort. Therefore, the agent also has incentive to exert effort at any time $t\leq\hat{t}$ in the dynamic model given contract $\hat{R}$. This implies that $\tau_{\hat{R}}\geq\tau_{R}$. Since the menu option $\hat{r}$ is chosen such that the agent’s incentive for exerting effort is binding at time $\tau_{R}$, the agent does not have incentive to exert effort at time $\tau_{R}+\Delta$ and hence $\tau_{\hat{R}}=\tau_{R}$. Finally, since $\hat{r}_{1}\leq r^{G}_{t,1}$ for any $t\leq\tau_{R}$, the agent’s expected reward for receiving good news signals at any time $t$ is weakly lower, and hence the ex ante reward constraint is satisfied given contract $\hat{R}$. ∎

### v3

@@ 6.1 Randomized Contracts
Our goal has been to determine the maximum effort implementable within the class of contracts defined in Section 2.2. As histories only include messages sent by the agent, we implicitly rule out the use of stochastic messages sent from the designer to the agent (as in, for instance, Deb et al., (2018)). Deterministic mechanisms have significant practical appeal, as it is not always obvious what a randomization might correspond to or how a mechanism designer could commit to implement this. These issues have been discussed extensively in the contracting literature; we refer the reader to discussions in Laffont and Martimort, (2002) as well as Bester and Strausz, (2001) on this point to avoid detours. Now, randomization provides no additional benefit in implementing maximum effort with probability 1. But it is natural to ask whether some designer objectives may yield benefits to randomization. Our analysis speaks to this question as well.
@@
Crucially, in randomized contracts, at any time $t$, the history of the randomization device $\left\{\varsigma_{t^{\prime}}\right\}_{t^{\prime}\leq t}$ is publicly revealed to the agent before determining his choice of effort or the message sent to the designer. The randomization revealed before $t$ affects the agent’s incentives after time $t$, and without it, such contracts reduce back to deterministic contracts. Implementing such randomized contracts would require either a public randomization device or the principal to commit to a random effort recommendation policy.
@@
At the same time, our analysis provides some insights as to why randomization can expand the set of implementable strategies. We previously observed that it may be possible to get the agent to work from $\mu_{1}$ to $\mu_{2}$ and to get the agent to work from $\mu_{2}$ to $\mu_{3}$, but not from $\mu_{1}$ to $\mu_{3}$. This would occur if the reward necessary to get the agent to work to $\mu_{3}$ were so high that the agent would “shirk-and-lie” at $\mu_{1}$. However, the agent may be willing to start working at $\mu_{1}$, not knowing whether the reward will be “high” or “low”—but once the agent starts working, the designer can randomly inflate or decrease the rewards of the agent. Once time has passed, the outcome of the randomization can be revealed, and if the rewards inflate, the agent has incentives to exert effort for longer absent a Poisson signal arrival—so that the realized stopping time increases for some realization of the randomization.
@@
We illustrate this intuition more formally and rigorously assuming perfect good-news learning, i.e., $\lambda_{1}^{G}>0$ and $\lambda_{0}^{G}=\lambda_{1}^{B}=\lambda_{0}^{B}=0$, where learning is perfect and only good news signal $G$ arrives with positive probability. We describe the resulting solution in Appendix OA 4. In this seting, our results imply that, under deterministic contracts, a V-shaped scoring rule with parameters $r_{0}=1$ and $r_{1}\in[0,1]$ implements the effort-maximizing contract $R$; in particular, Appendix OA 4 characterizes when in fact effort-maximizing contracts require $r_{1}<1$. Recall that we denote $\tau_{R}$ as the stopping time in the effort-maximizing deterministic contract and let $\mu^{N}_{\tau_{R}}$ be the stopping belief when no Poisson signal is observed. Let $\delta\in(0,1-r_{1}]$ be the maximum number such that (1) the agent has strict incentives to exert effort until time $\frac{2\tau_{R}}{3}$ absent signal arrival given menu options $(1,0)$ and $(0,r_{1}-\delta)$; and (2) the agent can be incentivized to exert effort given menu options $(1,0)$ and $(0,r_{1}+\delta)$ given belief $\mu^{N}_{\frac{\tau_{R}}{2}}$. Consider the randomized contract $\hat{R}$ that provides menu options $(1,0)$ and $(0,r_{1}-\delta)$ from time 00 to $\frac{\tau_{R}}{2}$, and after time $\frac{\tau_{R}}{2}$, offers menu options $(1,0)$ and $(0,r_{1}+\delta)$ with probability $\epsilon^{2}$, offers menu options $(1,0)$ and $(0,0)$ with probability $\epsilon$, and offers the same menu options $(1,0)$ and $(0,r_{1}-\delta)$ otherwise. With sufficiently small $\epsilon>0$, the agent still has incentives to exert effort at any time $t\leq\frac{\tau_{R}}{2}$. Moreover, after time $\frac{\tau_{R}}{2}$, with probability $\epsilon^{2}$, the realized menu options are $(1,0)$ and $(0,r_{1}+\delta)$, and the agent can be incentivized to exert effort to a time strictly larger than $\tau_{R}$ in the absence of a Poisson signal. [^14] This kind of modification may be of interest to a designer who only values extreme posterior beliefs. For instance, suppose the designer faced a decision problem where the possible decisions belonged to a set $A=\{0,1\}$. Consider a designer payoff function of $v(0,\theta)=0$ for all $\theta\in\Theta$, and $v(1,0)=1,v(1,1)=-\frac{1-\mu^{N}_{\tau_{R}}}{\mu^{N}_{\tau_{R}}}$; plainly, the designer only seeks to change her action from 1 to 0 if the posterior belief is below $\mu^{N}_{\tau_{R}}$. In this case, the payoff under any deterministic contract is 0. However, under the randomized contract outlined, the designer would obtain a positive payoff when implementing the identified randomized contract.
@@ 6.2 Non-invariant Environments
An important implication of extending our results to non-invariant environments is that they also apply to settings where the agent discounts future payoffs. Essentially, when the agents discounts future payoffs, it exactly corresponds to the the non-invariant environment with stationary information arrival but an exponentially decreasing cost function. Consequently, all our main results — such as the optimality of scoring rules in perfect-learning or single-signal environments — extend naturally to this setting.
6.3 Final Remarks
@@
We have articulated how dynamic rewards can expand the set of implementable strategies in a simple yet fundamentally dynamic information acquisition problem. The economic importance of contracting for information acquisition is self-evident, and most natural stories for why information acquisition is costly involve some dynamic element. Our goal has been to take such dynamics seriously, for learning technologies with a natural interpretation in terms of forecasters seeking a particular sought-after piece of falsifiable evidence, and under a class of contracts reflecting the provision of bonuses for correct advice.
@@
We have shown that whether the decision maker benefits from a contracting environment that facilitates dynamic reporting depends on the nature of the dynamic learning process. Along the way, we discussed how the relevant properties of the information acquisition technology have natural interpretations in various settings of practical interest. As our focus is on contracting under a general class of mechanisms, a fundamental difficulty underlying our exercise is the lack of any natural structure (e.g., stationarity) under an arbitrary dynamic contract. Such assumptions are often critical in similar settings. Despite this fundamental challenge, we provided simple, economically meaningful conditions such that maximum effort is implementable by a scoring rule and explained the extent to which these conditions are necessary for this conclusion to hold.
@@
There are many natural avenues for future work. Empirically, our results show how learning technologies influence the benefits of time variation to rewards. While we focus on a simple forecasting problem, such questions may be of interest when eliciting other kinds of information beyond a prediction of a future event—for example, if the learning process itself is privately known by the expert (Chambers and Lambert, , 2021). More broadly, we view questions regarding whether or not simple contracts are limited in power relative to dynamic mechanisms as a worthwhile agenda overall. Any further insights on these questions would prove valuable toward understanding how dynamics influence mechanism design for information acquisition, for both theory and practice.
@@ 2. Case 2:
Online Appendix for “Incentivizing Forecasters to Learn: Summarized vs. Unrestricted Advice”
@@
Appendix OA 2 Additional Review of Scoring Rules
@@
Appendix OA 3 Comparative Statics of Scoring Rules
@@
Appendix OA 4 Complete Solution under Perfect Good News Learning
OA 4.1 Effort-Maximizing Contracts
Here we describe our findings in the $\Delta\rightarrow 0$ limit, where a single signal reveals the state, i.e.,
$\lambda^{B}_{0}=\lambda^{B}_{1}=\lambda^{G}_{0}=0,~{}~{}\lambda^{G}_{1}>0.$
We take the horizon $T$ to be sufficiently large so that it will not be a binding constraint. Illustrating this solution highlights the tensions in maximizing the incentives to exert effort at different points in time.
For this learning environment, Theorem 2 implies that a scoring rule with two reward functions implements maximum effort: $(1,0)$, corresponding to a guess of state 0, and $(0,r_{1})$, corresponding to a guess of state 1. The value of $r_{1}$, the reward when guessing state 1 (correctly), depends on the initial prior, $D$ (which coincides with $\mu_{0}^{N}$).
Of course, when $\mu_{0}^{N}\in(\mu^{*},\mu^{**})$, if the agent works beliefs may eventually leave this region. If the agent had started at such a belief, more effort could be induced by setting $r_{1}=1$. On the other hand, it is clear why this reoptimization cannot help. When $\mu_{0}^{N}\in(\mu^{*},\mu^{**})$, $r_{1}<1$ will be set so that the initial moral hazard constraint binds—but if the agent expected $r_{1}$ to increase later, he would simply shirk and claim to have observed a Poisson signal once the reward is increased to $1$. Given that such adjustments are impossible, effort-maximizing mechanisms cannot utilize dynamics and are therefore static.
The lesson is that the tensions between optimizing incentives to exert effort both earlier and later may be unavoidable. A designer may be unable to re-optimize rewards because the re-optimized rewards would violate an incentive constraint at some other time. One deceptive aspect of this example is that the agent’s moral hazard constraint only ever binds at the stopping belief and (possibly) at their initial belief. A technical challenge we face in Section 4.3, for instance, is that if signals are not fully revealing, it may be that the moral hazard constraint binds somewhere “in between.” This feature will imply extra reward functions should be provided to the agent in the effort-maximizing scoring rule. Still, this example illustrates some of the intuition on how effort-maximizing rewards vary with the agent’s beliefs.
OA 4.2 Details Behind the Calculation
Here we provide some additional details behind the calculations in Section OA 4. We consider any scoring rule which involves the choice between $(r_{0},0)$ and $(0,r_{1})$, with the former being chosen when stopping in the absence of any signal arrival and the latter being chosen if one does occur. We note that stopping and accepting contract $(r_{0},0)$ delivers payoff $r_{0}(1-\mu_{t}^{N})$; if, at time $\tau$, the agent continues for a length $\Delta$ and then stops, the payoff is:
Imposing indifference between stopping and continuing yields:
From this, we obtain the following differential equation: